A MONO ATOMIC IDEAL GAS OF VOLUME 1 LITRE AT NTP Is ADIABITICALLY COMPRESSED TO HALF OF ITS OF ITS VOLUME .FIND THE CHANGE IN TEMPERAtURE CP/CV=5/3

Grade:11

1 Answers

nikhil
6 years ago
T2/ T1 = ( V1/ V2)^ ( n - 1)Now, V1/V2 = 2, At NTP , T1 = 20°C = 293 K.Therefore, T2 = 293* ( 2 )^[(5/3)-1] K